OpenStudy (anonymous):

please help step by step.. will fan & medal.. Find the limit of the function algebraically. limit as x approaches zero of quantity x squared plus two x divided by x to the fourth power.

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \frac{ x ^{2} + 2x }{ x ^{4}}\]

OpenStudy (misty1212):

HI!

OpenStudy (anonymous):

hello!

OpenStudy (misty1212):

factor and cancel first

OpenStudy (misty1212):

you get \[\lim_{x\to 0}\frac{x+2}{x^3}\]

OpenStudy (misty1212):

then numerator goes to 2, the denominator goes to zero, so no limit

OpenStudy (anonymous):

how does the numerator go to 2 and the demoninator go to zero?

OpenStudy (misty1212):

i will let you figure that out for yourself, since it is self evident replace x by zero

OpenStudy (anonymous):

i see it. can u show me how you factored this?

Nnesha (nnesha):

when you get 0 on the denominator by using direct substitution then first you should try to simplify the function. if you still get 0 after simplifying, that means limits doesn't exist but you still need t0 find limit from right and left when x approaches 2 to prove why limit does not exist

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @lxoser i see it. can u show me how you factored this? \(\color{blue}{\text{End of Quote}}\) \[\huge\rm \frac{ x^2+2x} { x^4 }\] factor the numerator there is x common so take out x from both terms x^2+2x =x(x+2) and then simplify

OpenStudy (kayders1997):

Have you learned Lhop rule yet? And differentiation?

OpenStudy (anonymous):

i havent..

OpenStudy (kayders1997):

Okay yes your going to have to factor and stuff than but first always Subsitituite you could get a quick answer from doing that in this case it doesn't work you will get 0/0 so you will have to factor

OpenStudy (kayders1997):

So what can you fActor out from the top?

OpenStudy (kayders1997):

What's something they both have in the top?

OpenStudy (anonymous):

x^2 + 2x = x(x+2)

OpenStudy (kayders1997):

Yes and than you can cancel another x from the bottom so what you have left is?

OpenStudy (kayders1997):

PS please write limit through each steps otherwise you could lose points

OpenStudy (kayders1997):

You okay?

OpenStudy (anonymous):

x + 2 / x^3

OpenStudy (kayders1997):

Yes now substitiute 0 for x because x is approaching 0